OpenStudy (x3_drummerchick):

how do i solve this? (will give medals!) 7^(x/2) = 3^(1 − x) i think im supposed to use substitution in some way

OpenStudy (x3_drummerchick):

\[7^{x/2}=3^{1-x}\]

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 7^{x/2}=3^{1-x} }\) \(\large\color{black}{ \displaystyle (\sqrt{7})^{x}=3^{1-x} }\) \(\large\color{black}{ \displaystyle (\sqrt{7})^{x}=3(3)^{-x} }\) \(\large\color{black}{ \displaystyle (\sqrt{7})^{x}=3\left(\frac{1}{3}\right)^{x} }\) \(\large\color{black}{ \displaystyle 3^x(\sqrt{7})^{x}=3 }\) \(\large\color{black}{ \displaystyle (3\sqrt{7})^{x}=3 }\) I don't want to do it entirely for you, so from here take it on please.

OpenStudy (mathstudent55):

Why not use logs? \(\Large 7^\frac{x}{2} = 3^{1 − x}\) Take the log of both sides: \(\large \log 7^\frac{x}{2} = \log 3^{1 − x}\) Use the exponent rule: \(\large \dfrac{x}{2} \log 7 = (1 - x) \log 3\) Multiply both sides by 2: \(\large x \log 7 = (2 - 2x) \log 3\) Now work on isolating x: \(\large x \log 7 = 2 \log 3 - 2x \log 3\) \(\large x \log 7 + 2x \log 3= 2 \log 3 \) \(\large x( \log 7 + 2 \log 3)= 2 \log 3 \) \(\large x = \dfrac{2 \log 3} {\log 7 + 2 \log 3}\)